# Deep Partially Linear Transformation Model for Right-Censored Survival Data

### v1

@@ 4 Simulation Studies
We first provide some computational details. The models are fitted by maximizing the log likelihood function (2) with respect to the parameters $\bm{\beta}$, $\widetilde{\gamma_{j}}$’s, $W_{k}$’s and $v_{k}$’s. For the DPLTM method, all the parameters are contained in one framework and simultaneously updated in each epoch, where the Adam optimizer (Kingma and Ba, 2014) is employed due to its efficiency and reliability. $\bm{\beta}$ and $\widetilde{\gamma_{j}}$’s are initialized to 0, while PyTorch’s default initialization algorithm is utilized for $W_{k}$’s and $v_{k}$’s.
The hyper-parameters in the simulations, including the number of hidden layers, the number of neurons in each hidden layer, the number of epochs, learning rate (Goodfellow, 2016), dropout rate (Srivastava et al., 2014) and the number of B-spline basis functions are tuned based on the log likelihood on the validation data via a grid search. We set the number of neurons in each hidden layer to be identical for convenience. We use cubic splines (i.e. $\ell$=3) to estimate $H_{0}(t)$ to achieve sufficient smoothness, with the number of interior knots $K_{n}$ chosen in the ranges $\lfloor n^{1/3}\rfloor,\cdots,2\lfloor n^{1/3}\rfloor$, and the number of basis functions $q_{n}=K_{n}+\ell+1$ can be determined. Candidates for other hyperparameters are presented in the Supplementary Material.
@@
The true transformation function $H_{0}(t)$ is set respectively to be $\log t$ for $r=0$, $\log(2e^{0.5t}-2)$ for $r=0.5$ and $\log(e^{t}-1)$ for $r=1$. Then we can generate survival time $U$ via its cumulative distribution function $F_{U}(t)=F_{\varepsilon}(H_{0}(t)+\bm{\beta}_{0}^{\top}\bm{Z}+g_{0}(\bm{X}))$ based on the inverse transform method. The censoring time is generated from a uniform distribution on $(0,c_{0})$, where the constant $c_{0}$ is chosen to approximately achieve the pre-specified censoring rate of 40% and 60% ($c_{0}=$2.95 or 0.85 for $r=0$, $c_{0}=$2.75 or 0.9 for $r=0.5$, $c_{0}=$2.55 or 1 for $r=1$, all kept the same for the three different cases of the underlying function $g_{0}(\bm{x})$).
We conduct 200 runs of simulation in each scenario with the sample size $n$=1000 and 2000. Our observations consist of $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$, where $T_{i}=\min\left\{U_{i},C_{i}\right\}$ and $\Delta_{i}=I(U_{i}\leq C_{i})$. We randomly split the samples into training data (80%) and validation data (20%), the former is used to fit models and obtain estimates, and the latter is utilized to tune the hyper-parameters. In addition, We generated $n_{test}=$200 and 400 test samples (corresponding to the cases of $n=$1000 and 2000 respectively) that are independent of the training samples for evaluation.
@@
$\text{RE}(\widehat{g})=\left\{\frac{\frac{1}{n_{test}}\sum_{i=1}^{n_{test}}\left[\left\{\widehat{g}(\bm{X}_{i})-\overline{\widehat{g}}\right\}-g_{0}(\bm{X}_{i})\right]^{2}}{\frac{1}{n_{test}}\sum_{i=1}^{n_{test}}\{g_{0}(\bm{X}_{i})\}^{2}}\right\}^{1/2},$
where $\overline{\widehat{g}}=\sum_{i=1}^{n_{test}}\widehat{g}(\bm{X}_{i})/n_{test}$.
We also evaluate and compare the predictive ability of the three methods using the concordance index (C-index). C-index is one of the most commonly used metrics to assess the predictive power of models in survival analysis. It measures the probability that the predicted survival times preserve the ranks of true survival times, which is defined as
@@
$\widehat{\text{C}}=\frac{\sum_{i=1}^{n_{test}}\sum_{j=1}^{n_{test}}\Delta_{i}1(T_{i}\leq T_{j})1(\widehat{\bm{\beta}}\bm{Z}_{i}+\widehat{g}(\bm{X}_{i})\geq\widehat{\bm{\beta}}\bm{Z}_{j}+\widehat{g}(\bm{X}_{j}))}{\sum_{i=1}^{n_{test}}\sum_{j=1}^{n_{test}}\Delta_{i}1(T_{i}\leq T_{j})}.$
The bias and standard deviation of the parametric estimate $\widehat{\bm{\beta}}$ derived from 200 simulation runs are presented in Table 1. It can be seen that the proposed DPLTM method provides asymptotically unbiased estimates in all scenarios. The biases for DPLTM are slightly higher than those for LTM and PLATM under Case 1, and PLATM under Case 2 respectively, which is expected because these two cases perfectly satisfy the simple linear and additive settings respectively. However, DPLTM greatly outperforms LTM and PLATM under Case 3 with a highly complicated true nonparametric function $g_{0}$, where the other two models are remarkably more biased than DPLTM and do not improve with increasing sample size. Moreover, the empirical standard deviation decreases steadily as $n$ increases for all three models under each simulation setting.
@@
In the Supplementary Material, we further measure the accuracy in estimating the transformation function $\widehat{H}$ with the weighted integrated squared error (WISE) and assess the calibration of the three methods using the integrated calibration index (ICI) (Austin et al., 2020). Similar conclusions can be drawn from results on these metrics. More importantly, we provide a comprehensive comparison between our method with the DPLCM method introduced by Zhong et al. (2022), and demonstrate that our method is generally more powerful and reliable than DPLCM due to its flexibility. In conclusion, these simulation results manifest that our DPLTM method performs consistently well with respect to both estimation and prediction in various settings.
Table 1: The bias and standard deviation of $\widehat{\bm{\beta}}$ for the DPLTM, LTM and PLATM methods.
$Z_{1}\sim N(0.5,0.5)$ $Z_{2}\sim Bernoulli(p=0.5)$ 40% censoring rate 60% censoring rate 40% censoring rate 60% censoring rate $r$ $n$ DPLTM LTM PLATM DPLTM LTM PLATM DPLTM LTM PLATM DPLTM LTM PLATM Case 1 0 1000 0.0318 0.0021 0.0335 0.0223 0.0059 0.0508 -0.0498 -0.0119 -0.0524 -0.0457 -0.0091 -0.0584 (Linear) (0.0961) (0.0874) (0.0997) (0.1144) (0.1075) (0.1260) (0.0959) (0.0868) (0.0948) (0.1224) (0.0983) (0.1094) 2000 0.0257 0.0014 0.0201 0.0031 0.0020 0.0245 -0.0287 -0.0029 -0.0230 -0.0172 -0.0037 -0.0289 (0.0678) (0.0601) (0.0614) (0.0813) (0.0716) (0.0757) (0.0721) (0.0625) (0.0662) (0.0862) (0.0779) (0.0816) 0.5 1000 0.0149 0.0034 0.0314 0.0164 0.0029 0.0490 -0.0390 -0.0101 -0.0523 -0.0429 -0.0077 -0.0599 (0.1189) (0.1133) (0.1265) (0.1375) (0.1347) (0.1513) (0.1234) (0.1117) (0.1186) (0.1461) (0.1216) (0.1320) 2000 0.0101 0.0012 0.0205 -0.0083 0.0022 0.0244 -0.0165 -0.0038 -0.0239 -0.0111 -0.0066 -0.0318 (0.0851) (0.0785) (0.0792) (0.1006) (0.0885) (0.0914) (0.0918) (0.0804) (0.0856) (0.1083) (0.0947) (0.0991) 1 1000 -0.0132 -0.0029 0.0306 -0.0222 0.0026 0.0414 -0.0247 -0.0135 -0.0542 -0.0317 -0.0102 -0.0623 (0.1354) (0.1400) (0.1528) (0.1541) (0.1515) (0.1721) (0.1289) (0.1383) (0.1449) (0.1530) (0.1444) (0.1545) 2000 -0.0042 0.0015 0.0172 -0.0089 0.0020 0.0237 -0.0112 -0.0051 -0.0245 -0.0114 -0.0065 -0.0329 (0.0954) (0.0944) (0.0995) (0.1030) (0.1004) (0.1046) (0.1000) (0.0952) (0.1006) (0.1158) (0.1096) (0.1144) Case 2 0 1000 -0.0511 -0.3844 0.0322 -0.0712 -0.3262 0.0554 0.0463 0.3770 -0.0592 0.0365 0.3189 -0.0537 (Additive) (0.1068) (0.0907) (0.0968) (0.1288) (0.1126) (0.1231) (0.1090) (0.0868) (0.0931) (0.1298) (0.1002) (0.1155) 2000 -0.0321 -0.3872 0.0122 -0.0523 -0.3284 0.0221 -0.0226 0.3816 -0.0165 0.0184 0.3271 -0.0248 (0.0684) (0.0553) (0.0596) (0.0850) (0.0672) (0.0742) (0.0688) (0.0643) (0.0634) (0.0816) (0.0768) (0.0773) 0.5 1000 -0.0479 -0.2892 0.0340 -0.0564 -0.2638 0.0570 0.0372 0.2767 -0.0626 0.0326 0.2546 -0.0572 (0.1311) (0.1134) (0.1229) (0.1521) (0.1350) (0.1526) (0.1374) (0.1088) (0.1145) (0.1616) (0.1241) (0.1513) 2000 -0.0241 -0.2895 0.0134 -0.0383 -0.2657 0.0243 -0.0211 0.2830 -0.0190 -0.0178 0.2630 -0.0283 (0.0891) (0.0718) (0.0790) (0.1008) (0.0770) (0.0923) (0.0868) (0.0815) (0.0824) (0.0987) (0.0914) (0.0972) 1 1000 -0.0426 -0.2296 0.0292 -0.0610 -0.2226 0.0552 0.0307 0.2231 -0.0672 0.0299 0.2125 -0.0615 (0.1385) (0.1323) (0.1472) (0.1635) (0.1480) (0.1734) (0.1495) (0.1284) (0.1395) (0.1713) (0.1353) (0.1691) 2000 -0.0314 -0.2278 0.0145 -0.0486 -0.2309 0.0220 -0.0112 0.2231 -0.0182 -0.0102 0.2237 -0.0278 (0.1012) (0.1350) (0.0981) (0.1127) (0.0883) (0.1082) (0.0941) (0.1241) (0.0998) (0.1073) (0.1039) (0.1151) Case 3 0 1000 -0.0499 -0.4616 -0.3734 -0.0613 -0.4379 -0.3500 0.0695 0.4433 0.3718 0.0512 0.4030 0.3432 (Deep) (0.1106) (0.0844) (0.1097) (0.1284) (0.1017) (0.1267) (0.1127) (0.0864) (0.1081) (0.1369) (0.1097) (0.1218) 2000 -0.0333 -0.4558 -0.3941 -0.0471 -0.4275 -0.3650 0.0361 0.4648 0.4085 0.0385 0.4305 0.3860 (0.0834) (0.0608) (0.0669) (0.0856) (0.0738) (0.0810) (0.0771) (0.0575) (0.0662) (0.0824) (0.0715) (0.0786) 0.5 1000 -0.0601 -0.3668 -0.2926 -0.0643 -0.3761 -0.2899 0.0606 0.3426 0.2876 0.0667 0.3352 0.2827 (0.1138) (0.1053) (0.1264) (0,1478) (0.1167) (0.1457) (0.1212) (0.1072) (0.1351) (0.1317) (0.1262) (0.1377) 2000 -0.0280 -0.3599 -0.3061 -0.0483 -0.3599 -0.3031 0.0295 0.3669 0.3215 0.0487 0.3640 0.3242 (0.0893) (0.0736) (0.0785) (0.0980) (0.0852) (0.0915) (0.0881) (0.0704) (0.0817) (0.0935) (0.0856) (0.0900) 1 1000 -0.0594 -0.2996 -0.2263 -0.0595 -0.3294 -0.2454 0.0519 0.2717 0.2231 0.0661 0.2842 0.2379 (0.1456) (0.1267) (0.1429) (0.1473) (0.1322) (0.1639) (0.1350) (0.1291) (0.1546) (0.1611) (0.1426) (0.1567) 2000 -0.0449 -0.2913 -0.2445 -0.0441 -0.3093 -0.2590 0.0371 0.2987 0.2603 0.0469 0.3157 0.2796 (0.1071) (0.0887) (0.0918) (0.0939) (0.0998) (0.0984) (0.1040) (0.0834) (0.0976) (0.1082) (0.0947) (0.1039)
Table 2: Empirical coverage probability of 95% confidence intervals for $\bm{\beta}_{0}$ for the DPLTM, LTM and PLATM methods.
$Z_{1}\sim N(0.5,0.5)$ $Z_{2}\sim Bernoulli(p=0.5)$ 40% censoring rate 60% censoring rate 40% censoring rate 60% censoring rate $r$ $n$ DPLTM LTM PLATM DPLTM LTM PLATM DPLTM LTM PLATM DPLTM LTM PLATM Case 1 0 1000 0.950 0.950 0.925 0.960 0.945 0.940 0.945 0.965 0.935 0.965 0.960 0.920 (Linear) 2000 0.955 0.930 0.935 0.950 0.950 0.935 0.955 0.960 0.945 0.950 0.955 0.930 0.5 1000 0.945 0.960 0.945 0.965 0.945 0.940 0.970 0.970 0.930 0.950 0.975 0.930 2000 0.955 0.940 0.925 0.940 0.960 0.935 0.960 0.960 0.945 0.950 0.960 0.935 1 1000 0.950 0.960 0.935 0.950 0.960 0.925 0.945 0.970 0.930 0.945 0.970 0.915 2000 0.940 0.935 0.930 0.960 0.960 0.950 0.975 0.955 0.945 0.945 0.970 0.930 Case 2 0 1000 0.935 0.020 0.920 0.925 0.140 0.915 0.940 0.010 0.935 0.940 0.125 0.910 (Additive) 2000 0.945 0.000 0.940 0.930 0.000 0.930 0.940 0.000 0.920 0.930 0.005 0.925 0.5 1000 0.920 0.235 0.930 0.935 0.410 0.920 0.955 0.275 0.910 0.930 0.475 0.910 2000 0.945 0.030 0.940 0.950 0.130 0.945 0.950 0.040 0.930 0.940 0.155 0.945 1 1000 0.925 0.575 0.940 0.940 0.625 0.930 0.930 0.615 0.920 0.935 0.675 0.930 2000 0.945 0.285 0.950 0.935 0.340 0.960 0.940 0.350 0.940 0.955 0.385 0.935 Case 3 0 1000 0.925 0.000 0.045 0.910 0.005 0.150 0.925 0.000 0.045 0.915 0.045 0.170 (Deep) 2000 0.945 0.000 0.000 0.920 0.000 0.000 0.920 0.000 0.000 0.935 0.000 0.010 0.5 1000 0.930 0.075 0.255 0.915 0.125 0.415 0.940 0.095 0.310 0.910 0.215 0.405 2000 0.910 0.010 0.030 0.925 0.015 0.080 0.935 0.000 0.020 0.920 0.015 0.055 1 1000 0.910 0.360 0.635 0.915 0.355 0.655 0.920 0.415 0.590 0.905 0.460 0.590 2000 0.925 0.105 0.285 0.930 0.125 0.290 0.945 0.090 0.245 0.920 0.105 0.200
Table 3: The average and standard deviation of the relative error of $\widehat{g}$ for the DPLTM, LTM and PLATM methods.
      40% censoring rate       60% censoring rate       $r$       $n$       DPLTM       LTM       PLATM       DPLTM       LTM       PLATM       Case 1       0       1000       0.1164       0.0588       0.0860       0.0864       0.0598       0.1243       (Linear)       (0.0248)       (0.0403)       (0.0293)       (0.0302)       (0.0396)       (0.0219)       2000       0.0796       0.0440       0.0546       0.0854       0.0443       0.0639       (0.0191)       (0.0300)       (0.0249)       (0.0179)       (0.0298)       (0.0170)       0.5       1000       0.1645       0.0597       0.0898       0.1135       0.0607       0.1270       (0.0199)       (0.0397)       (0.0283)       (0.0252)       (0.0391)       (0.0215)       2000       0.1105       0.0442       0.0577       0.0882       0.0446       0.0674       (0.0156)       (0.0298)       (0.0239)       (0.0177)       (0.0296)       (0.0161)       1       1000       0.1516       0.0591       0.0818       0.1369       0.0609       0.1077       (0.0212)       (0.0402)       (0.0306)       (0.0232)       (0.0390)       (0.0245)       2000       0.1156       0.0445       0.0543       0.0996       0.0451       0.0641       (0.0151)       (0.0297)       (0.0250)       (0.0163)       (0.0292)       (0.0180)       Case 2       0       1000       0.3301       0.7841       0.0857       0.3491       0.7721       0.0945       (Additive)       (0.0238)       (0.0221)       (0.0326)       (0.0303)       (0.0248)       (0.0303)       2000       0.2871       0.7845       0.0526       0.2794       0.7729       0.0570       (0.0111)       (0.0160)       (0.0272)       (0.0154)       (0.0179)       (0.0257)       0.5       1000       0.3511       0.7693       0.0878       0.3609       0.7647       0.0999       (0.0337)       (0.0253)       (0.0321)       (0.0346)       (0.0267)       (0.0292)       2000       0.2827       0.7700       0.0540       0.3007       0.7653       0.0592       (0.0132)       (0.0182)       (0.0267)       (0.0115)       (0.0193)       (0.0251)       1       1000       0.3075       0.7641       0.0834       0.3358       0.7621       0.0992       (0.0320)       (0.0266)       (0.0333)       (0.0340)       (0.0275)       (0.0293)       2000       0.2629       0.7644       0.0528       0.3136       0.7628       0.0585       (0.0130)       (0.0192)       (0.0271)       (0.0120)       (0.0198)       (0.0253)       Case 3       0       1000       0.4069       0.9158       0.8524       0.4835       0.9086       0.8444       (Deep)       (0.0761)       (0.0140)       (0.0176)       (0.0748)       (0.0161)       (0.0212)       2000       0.3621       0.9172       0.8574       0.4390       0.9100       0.8490       (0.0404)       (0.0097)       (0.0124)       (0.0417)       (0.0112)       (0.0151)       0.5       1000       0.4716       0.9059       0.8415       0.4739       0.9032       0.8399       (0.0836)       (0.0170)       (0.0204)       (0.0872)       (0.0180)       (0.0228)       2000       0.3887       0.9075       0.8462       0.4186       0.9045       0.8438       (0.0565)       (0.0119)       (0.0147)       (0.0579)       (0.0126)       (0.0166)       1       1000       0.4032       0.9011       0.8375       0.4297       0.9002       0.8366       (0.0933)       (0.0187)       (0.0216)       (0.0812)       (0.0192)       (0.0230)       2000       0.3390       0.9026       0.8417       0.3627       0.9017       0.8405       (0.0598)       (0.0134)       (0.0159)       (0.0523)       (0.0135)       (0.0169)
Table 4: The average and standard deviation of the C-index for the DPLTM, LTM and PLATM methods.
      40% censoring rate       60% censoring rate       $r$       $n$       DPLTM       LTM       PLATM       DPLTM       LTM       PLATM       Case 1       0       1000       0.8374       0.8379       0.8298       0.8474       0.8475       0.8402       (Linear)       (0.0171)       (0.0167)       (0.0172)       (0.0208)       (0.0201)       (0.0209)       2000       0.8358       0.8375       0.8334       0.8461       0.8484       0.8448       (0.0121)       (0.0112)       (0.0113)       (0.0140)       (0.0134)       (0.0137)       0.5       1000       0.8153       0.8162       0.8064       0.8281       0.8292       0.8196       (0.0195)       (0.0184)       (0.0189)       (0.0229)       (0.0217)       (0.0225)       2000       0.8155       0.8148       0.8098       0.8221       0.8299       0.8246       (0.0139)       (0.0123)       (0.0126)       (0.0152)       (0.0143)       (0.0146)       1       1000       0.8067       0.8042       0.8106       0.8058       0.8110       0.8198       (0.0192)       (0.0199)       (0.0200)       (0.0228)       (0.0233)       (0.0239)       2000       0.8161       0.8020       0.8062       0.8063       0.8105       0.8154       (0.0140)       (0.0129)       (0.0130)       (0.0153)       (0.0151)       (0.0154)       Case 2       0       1000       0.8192       0.7265       0.8324       0.8239       0.7462       0.8441       (Additive)       (0.0183)       (0.0207)       (0.0167)       (0.0224)       (0.0248)       (0.0190)       2000       0.8226       0.7269       0.8349       0.8287       0.7467       0.8462       (0.0123)       (0.0163)       (0.0126)       (0.0146)       (0.0194)       (0.0151)       0.5       1000       0.7990       0.7192       0.8118       0.7992       0.7360       0.8254       (0.0218)       (0.0221)       (0.0176)       (0.0249)       (0.0262)       (0.0203)       2000       0.8037       0.7188       0.8124       0.8146       0.7358       0.8272       (0.0137)       (0.0170)       (0.0137)       (0.0152)       (0.0202)       (0.0162)       1       1000       0.7845       0.6981       0.7883       0.7928       0.7183       0.8070       (0.0214)       (0.0214)       (0.0186)       (0.0250)       (0.0253)       (0.0213)       2000       0.7946       0.6975       0.7898       0.8121       0.7184       0.8086       (0.0139)       (0.0160)       (0.0146)       (0.0162)       (0.0197)       (0.0170)       Case 3       0       1000       0.7814       0.6743       0.7022       0.7830       0.6927       0.7219       (Deep)       (0.0235)       (0.0246)       (0.0244)       (0.0304)       (0.0284)       (0.0271)       2000       0.7904       0.6757       0.7030       0.7936       0.6958       0.7227       (0.0147)       (0.0168)       (0.0165)       (0.0170)       (0.0198)       (0.0183)       0.5       1000       0.7786       0.6696       0.6927       0.7820       0.6852       0.7107       (0.0237)       (0.0258)       (0.0246)       (0.0280)       (0.0294)       (0.0282)       2000       0.7809       0.6711       0.6935       0.8030       0.6857       0.7119       (0.0171)       (0.0180)       (0.0169)       (0.0201)       (0.0205)       (0.0192)       1       1000       0.7887       0.6538       0.6805       0.7965       0.6710       0.6985       (0.0236)       (0.0235)       (0.0252)       (0.0294)       (0.0293)       (0.0285)       2000       0.7904       0.6549       0.6816       0.8139       0.6722       0.7004       (0.0166)       (0.0169)       (0.0171)       (0.0197)       (0.0201)       (0.0193)

### v2

@@ 4 Simulation Studies
We first provide some computational details. The models are fitted by maximizing the log likelihood function (2) with respect to the parameters $\bm{\beta}$, $\widetilde{\gamma}_{j}$’s, $W_{k}$’s and $v_{k}$’s. For the DPLTM method, all the parameters are contained in one framework and simultaneously updated in each epoch, where the Adam optimizer (Kingma and Ba, 2014) is employed due to its efficiency and reliability. All components of $\bm{\beta}$ and all $\widetilde{\gamma}_{j}$’s are initialized to 0 and -1, respectively, while PyTorch’s default initialization algorithm is utilized for $W_{k}$’s and $v_{k}$’s.
The hyper-parameters in the simulations, including the number of hidden layers, the number of neurons in each hidden layer, the number of epochs, learning rate (Goodfellow, 2016), dropout rate (Srivastava et al., 2014) and the number of B-spline basis functions are tuned based on the log likelihood on the validation data via a grid search. We set the number of neurons in each hidden layer to be identical for convenience. We use cubic splines (i.e. $\ell$=3) to estimate $H_{0}(t)$ to achieve sufficient smoothness, with the number of interior knots $K_{n}$ chosen in the ranges $\lfloor n^{1/3}\rfloor,\cdots,2\lfloor n^{1/3}\rfloor$, and the number of basis functions $q_{n}=K_{n}+\ell+1$ can be determined. Candidates for other hyperparameters are presented in the Supplementary Materials.
@@
The true transformation function $H_{0}(t)$ is set respectively to be $\log t$ for $r=0$, $\log(2e^{0.5t}-2)$ for $r=0.5$ and $\log(e^{t}-1)$ for $r=1$. Then we can generate survival time $U$ via its distribution function $F_{U}(t)=F_{\varepsilon}(H_{0}(t)+\bm{\beta}_{0}^{\top}\bm{Z}+g_{0}(\bm{X}))$ based on the inverse transform method. The censoring time is generated from a uniform distribution on $(0,c_{0})$, where the constant $c_{0}$ is chosen to approximately achieve the pre-specified censoring rate of 40% and 60% ($c_{0}=$2.95 or 0.85 for $r=0$, $c_{0}=$2.75 or 0.9 for $r=0.5$, $c_{0}=$2.55 or 1 for $r=1$, all kept the same for the three different cases of the underlying function $g_{0}(\bm{x})$).
We conduct 200 runs of simulation in each scenario with the sample size $n$=1000 and 2000. Our observations consist of $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$, where $T_{i}=\min\left\{U_{i},C_{i}\right\}$ and $\Delta_{i}=I(U_{i}\leq C_{i})$. We randomly split the samples into training data (80%) and validation data (20%), the former is used to fit models and obtain estimates, and the latter is utilized to tune the hyper-parameters. In addition, We generated $n_{\text{test}}=$200 and 400 test samples (corresponding to the cases of $n=$1000 and 2000 respectively) that are independent of the training samples for evaluation.
@@
$\text{RE}(\widehat{g})=\left\{\frac{\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\left[\left\{\widehat{g}(\bm{X}_{i})-\overline{\widehat{g}}\right\}-g_{0}(\bm{X}_{i})\right]^{2}}{\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\{g_{0}(\bm{X}_{i})\}^{2}}\right\}^{1/2},$
where $\overline{\widehat{g}}=\sum_{i=1}^{n_{\text{test}}}\widehat{g}(\bm{X}_{i})/n_{\text{test}}$.
We also evaluate and compare the predictive ability of the three methods using the concordance index (C-index) by Harrell et al. (1982). C-index is one of the most commonly used metrics to assess the predictive power of models in survival analysis. It measures the probability that the predicted survival times preserve the ranks of true survival times, which is defined as
@@
$\widehat{\text{C}}=\frac{\sum_{i=1}^{n_{\text{test}}}\sum_{j=1}^{n_{\text{test}}}\Delta_{i}1(T_{i}\leq T_{j})1(\widehat{\bm{\beta}}\bm{Z}_{i}+\widehat{g}(\bm{X}_{i})\geq\widehat{\bm{\beta}}\bm{Z}_{j}+\widehat{g}(\bm{X}_{j}))}{\sum_{i=1}^{n_{\text{test}}}\sum_{j=1}^{n_{\text{test}}}\Delta_{i}1(T_{i}\leq T_{j})}.$
The bias and standard deviation of the parametric estimate $\widehat{\bm{\beta}}$ derived from 200 simulation runs are presented in Table 1. It can be seen that the proposed DPLTM method provides asymptotically unbiased estimates in all scenarios. The biases for DPLTM are sometimes slightly higher than those for LTM and PLATM under Case 1, and PLATM under Case 2 respectively, which is expected because these two cases perfectly satisfy the simple linear and additive settings respectively. However, DPLTM greatly outperforms LTM and PLATM under Case 3 with a highly complicated true nonparametric function $g_{0}$, where the other two models are remarkably more biased than DPLTM and do not improve with increasing sample size. Moreover, the empirical standard deviation decreases steadily as $n$ increases for all three models under each simulation setting.
@@
In the Supplementary Materials, we further measure the accuracy in estimating the transformation function $\widehat{H}$ with the weighted integrated squared error (WISE) and assess the calibration of the three methods using the integrated calibration index (ICI) (Austin et al., 2020). Similar conclusions can be drawn from results on these metrics. More importantly, we provide a comprehensive comparison between our method with the DPLCM method introduced by Zhong et al. (2022), and demonstrate that our method is generally more powerful and reliable than DPLCM due to its flexibility. In conclusion, these simulation results manifest that our DPLTM method performs consistently well with respect to both estimation and prediction in various settings.